OpenStudy (anonymous):

The ordered pair below is a solution to which equation? (40, 5) A. y = 8x B. y = 5x C. y = 1/8x D. y = 1/5x

OpenStudy (anonymous):

C Plug in 40 for the x value. So you have 1/8 * 40/1 = 40/8 \[40\div8= 5\]

OpenStudy (anonymous):

Find the equation of the line .

OpenStudy (anonymous):

Find the slope by dividing the raise over the run. m= (12-4)/(3-1) = 4 Take one of the points and plug into the slope intercept form y= mx + b For point (1,4) you get 4 = 4(1) + b So b = 0 So the equation of the line is y = 4x + 0